OpenStudy (anonymous):

solve the system by using elimination or elimination with multiplication. Write the solution as an ordered pair. (x'y) x+y=16 x=y=-2

OpenStudy (anonymous):

Solve the system by using elimination or elimination with multiplication. x + y = 16 and -1(x + y = -2) x+y=16 -x-y=2 add these together, both x and y will cancel out, so 0=18 therefore no solution

OpenStudy (anonymous):

need anything else?